OpenStudy (anonymous):

how do u solve 2x divided by 5 =14

OpenStudy (anonymous):

\[\frac{ 2x }{ 5 }=14\]

OpenStudy (anonymous):

what you need to do is isolate x, to do this you must "undo" everything that is being done to it

OpenStudy (anonymous):

and do it to the other side

OpenStudy (anonymous):

so first of all, 2x is being divided by 5, to undo this division, we can multiply it by 5

OpenStudy (anonymous):

but everything we do to one side of an equation we need to do to the other side so that they remain equal

OpenStudy (anonymous):

so we multiply 14 by 5 as well to get\[5 \times \frac{ 2x }{ 5} = 5 \times 14\]

OpenStudy (anonymous):

the fives on the left side cancel and we can simplify the expression 5 x 14 to get 70, and now the equation is \[2x = 70\]

OpenStudy (anonymous):

but x is not isolated, so we need to "undo" what is being done to it.

OpenStudy (anonymous):

x is being multiplied by 2 so to undo that we divide by 2

OpenStudy (anonymous):

\[\frac{ 2x }{ 2 } = \frac{ 70 }{ 2 }\]

OpenStudy (anonymous):

the twos on the left cancel, and we can simplify the expression on the right to 35 leaving us with x = 35

OpenStudy (anonymous):

thank you very much